a system of grating has two wavelength 5.0 x 10^-7 and 5.2 x 10^-7m. the focal length is 2.0n and the grating separation is 2.0×10^-6m find the change in linear separatiom per unit increase in the specttal order of two wavelength
1
Expert's answer
2021-10-07T09:45:36-0400

Write down the conditions for maxima and determine the angle:


"d\sin\theta_1=m\lambda_1,\\
d\sin\theta_2=m\lambda_2.\\\space\\
\theta_1=\arcsin\frac{m\lambda_1}{d},\\\space\\
\theta_2=\arcsin\frac{m\lambda_2}{d}."

For the length between the grating and screen 2.0 m, find the change in linear separation:


"\Delta z(m)=D(\tan\theta_2-\tan\theta_1)=\\\space\\
=D(\tan\arcsin\frac{m\lambda_2}{d}-\tan\arcsin\frac{m\lambda_1}{d})=\\\space\\
=2[\tan\arcsin(0.26m)-\tan\arcsin(0.25m)]."